# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v2

@@ C.4 Zero-shot Forecasting full results
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, cells = c, colspec = Q[5em,c]Q[1.5em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c], cell11,3,5,7,9,11,13,15,17 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ C.5 Classification Full Results
{tblr}
colsep = 4pt, rowsep = 1pt, cell11 = r=2, cell12 = c=2, cell14 = c=2, cell16 = r=2, cell17 = c=9, cell116 = c=2, cell118 = r=2, cell119 = c=4, cells = c,m, hline1,14 = 1-220.15em, hline3,13 = 1-220.05em, vline2,4,6,7,16,18,19 = 1abovepos = -1.2, vline2,4,6,7,16,18,19 = 2belowpos = -1.2, vline2,4,6,7,16,18,19 = 4-11, vline2,4,6,7,16,18,19 = 3abovepos = -1.2, vline2,4,6,7,16,18,19 = 12belowpos = -1.2, vline2,4,6,7,16,18,19 = 13abovepos = -1.2, belowpos = -1.2, Methods & Classical methods RNN TCN Transformers MLP TimesNet LLM-based
@@ Appendix E Schematic Diagram of the Framework
Firstly, for the fine-grained branch, consider the scenario in which the input TS embeddings are segmented into 2 subsequences, each comprising four embeddings. Thus, $TS_{sub}^{1}$ is $[{\bm{e}}_{1,1},{\bm{e}}_{1,2},{\bm{e}}_{1,3},{\bm{e}}_{1,4}]$, and $TS_{sub}^{2}$ is $[{\bm{e}}_{2,1},{\bm{e}}_{2,2},{\bm{e}}_{2,3},{\bm{e}}_{2,4}]$, where ${\bm{e}}_{i,j}$ indicates $j$-th embedding in subsequence $i$. Similarly, ${\bm{z}}_{i,j}$ refers to $j$-th embedding in the prompt of subsequence $i$. Here, $[{\bm{z}}_{1,1},{\bm{z}}_{1,2}]=[{\bm{z}}_{2,1},{\bm{z}}_{2,2}]$. Ultimately, Eq.5 is instantiated as:
$[{\bm{e}}_{1,1},{\bm{e}}_{1,2},{\bm{e}}_{1,3},{\bm{e}}_{1,4},{\bm{z}}_{1,1},{\bm{z}}_{1,2},{\bm{e}}_{2,1},{\bm{e}}_{2,2},{\bm{e}}_{2,3},{\bm{e}}_{2,4},{\bm{z}}_{2,1},{\bm{z}}_{2,2}].$
@@
Thirdly, for the coarse-grained branch, it is essential to inform the LLM that a time series should be treated as a whole. Thus, $TS_{sub}^{i}$ must be mapped to individual node embedding by a linear layer. To align the scales, the prompt embeddings are also mapped to a node embedding. Thus, the coarse-grained sequences can be denoted as $[\tilde{\bm{e}}_{1},\tilde{\mathbf{z}}^{(1)},\tilde{\bm{e}}_{2},\tilde{\mathbf{z}}^{(2)}]$ (instantiation of Eq.6). Additionally, the graph construction logic is consistent with that of the fine-grained branch.
@@ F.2 Scale analysis
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, colspec = Q[75]Q[75]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87], cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell21 = c=2, cell31 = r=5, cell81 = r=5, hline1,13 = 1-100.15em, hline2,3,8 = 1-100.05em, vline3,4,5,6,7,8,9,10 = 1,2abovepos = -1.2, belowpos = -1.2, vline3,4,5,6,7,8,9,10 = 4-11, vline2 = 4-6, vline2 = 3abovepos = -1.2, vline2 = 7belowpos = -1.2, vline2 = 9-11, vline2 = 8abovepos = -1.2, vline2 = 12belowpos = -1.2, vline3,4,5,6,7,8,9,10 = 3abovepos = -1.2, vline3,4,5,6,7,8,9,10 = 12belowpos = -1.2, Examples amount & 1 2 3 4
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, colspec = Q[100]Q[100]Q[100], hline1,8 = 1-30.15em, hline2 = 1-30.05em, vline2,3 = 1abovepos = -1.2, belowpos = -1.2, vline2,3 = 2abovepos = -1.2, vline2,3 = 7belowpos = -1.2, vline2,3 = 3-6, Training data ratios & ETTh1 ETTm1
@@ F.3 Fully tuning analysis
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, colspec = Q[250]Q[100]Q[100], hline1,6 = 1-30.15em, hline2,4 = 1-30.05em, vline2,3 = 1abovepos = -1.2, belowpos = -1.2, vline2,3 = 2,4abovepos = -1.2, vline2,3 = 3,5belowpos = -1.2, Method & ETTh1 ETTm1
@@ F.4 Comparison between VCA and other LLM-based methods
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = colspec = Q[117]Q[81]Q[81]Q[81]Q[81]Q[81]Q[81]Q[81]Q[81]Q[81]Q[81], cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, hline1 = 1-110.15em, hline2,3,11 = 1-110.05em, hline12 = 1-110.15em, Methods & FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, cell11 = c=2, cell21 = r=3, hline1 = 1-70.15em, hline2 = 1-70.05em, hline5 = 1-70.15em, Methods & FSCA VCA S2IP-LLM Time-LLM GPT4TS
Average SMAPE 11.828 11.889 12.021 12.494 12.690
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, hline1 = 1-110.15em, hline2 = 1-110.05em, hline3 = 1-110.05em, hline7 = 1-110.05em, hline8 = 1-110.15em, Methods & FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, hline1 = 1-110.15em, hline2 = 1-110.05em, hline3 = 1-110.05em, hline11 = 1-110.05em, hline12 = 1-110.15em, Methods & FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, cell11 = r=2, cell12 = c=5, hline1 = 1-60.15em, hline2 = 2-60.05em, hline3 = 1-60.05em, hline13 = 1-60.05em, hline14 = 1-60.15em, Methods & LLM-based
@@ F.5 Experimental efficiency analysis
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, width = cells = c, hline1,8 = 1-60.15em, hline2,5 = 1-60.05em, & Training Params  Training Params Percentages  Training Time for 1 iteration(s)  Inference Time for 1 iteration(s)
@@ F.6 Comparison of different prompt types
{tblr}
colsep = 2.2pt, rowsep = 1.pt, width = cells = c, hline1, 5 = 1-50.15em, hline2 = 1-110.05em, Variant & ETTh1 ETTm1 ETTm1 ETTm2
@@ F.7 Comparison of different network types implementing Context-Alignment
{tblr}
colsep = 2.2pt, rowsep = 1.pt, width = cells = c, hline1,7 = 1-50.15em, hline2,4 = 1-110.05em, Variant & ETTh1 ETTm1 ETTm1 ETTm2
@@ F.8 Comparison with TS Foundation Models
{tblr}
colsep = 2.2pt, rowsep = 1.pt, width = cells = c, cell11 = r=2, cell12 = c=4, cell16 = c=2, hline1,14 = 1-70.15em, hline3 = 1-70.05em, hline2 = 2-70.05em, vline2,6 = 4-12, vline2,6 = 3abovepos = -1.2, vline2,6 = 13belowpos = -1.2, vline6 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & LLM-based Models TS Foundation Models
@@
{tblr}
colsep = 2.2pt, rowsep = 1.pt, width = cells = c, cell11 = r=2, cell12 = c=8, cell110 = c=6, cell22 = c=2, cell24 = c=2, cell26 = c=2, cell28 = c=2, cell210 = c=2, cell212 = c=2, cell214 = c=2, hline1,12 = 1-150.15em, hline3,4 = 1-150.05em, hline2 = 2-150.05em, vline2,10 = 5-10, vline10 = 4abovepos = -1.2, vline10 = 11belowpos = -1.2, vline10 = 1,2,3abovepos = -1.2, belowpos = -1.2, vline2 = 4abovepos = -1.2, vline2 = 11belowpos = -1.2, Methods & LLM-based Models TS Foundation Models

### v3

@@ C.4 Zero-shot Forecasting full results
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ C.5 Classification Full Results
Methods Classical methods RNN TCN Transformers MLP TimesNet LLM-based
@@ Appendix E Schematic Diagram of the Framework
Firstly, for the fine-grained branch, consider the scenario in which the input TS embeddings are segmented into 2 subsequences, each comprising four embeddings. Thus, $TS_{sub}^{1}$ is ${\bm{e}}_{1,1},{\bm{e}}_{1,2},{\bm{e}}_{1,3},{\bm{e}}_{1,4}]$, and $TS_{sub}^{2}$ is ${\bm{e}}_{2,1},{\bm{e}}_{2,2},{\bm{e}}_{2,3},{\bm{e}}_{2,4}]$, where ${\bm{e}}_{i,j}$ indicates $j$-th embedding in subsequence $i$. Similarly, ${\bm{z}}_{i,j}$ refers to $j$-th embedding in the prompt of subsequence $i$. Here, ${\bm{z}}_{1,1},{\bm{z}}_{1,2}]=[{\bm{z}}_{2,1},{\bm{z}}_{2,2}]$. Ultimately, Eq.5 is instantiated as:
${\bm{e}}_{1,1},{\bm{e}}_{1,2},{\bm{e}}_{1,3},{\bm{e}}_{1,4},{\bm{z}}_{1,1},{\bm{z}}_{1,2},{\bm{e}}_{2,1},{\bm{e}}_{2,2},{\bm{e}}_{2,3},{\bm{e}}_{2,4},{\bm{z}}_{2,1},{\bm{z}}_{2,2}].$
@@
Thirdly, for the coarse-grained branch, it is essential to inform the LLM that a time series should be treated as a whole. Thus, $TS_{sub}^{i}$ must be mapped to individual node embedding by a linear layer. To align the scales, the prompt embeddings are also mapped to a node embedding. Thus, the coarse-grained sequences can be denoted as $\tilde{\bm{e}}_{1},\tilde{\mathbf{z}}^{(1)},\tilde{\bm{e}}_{2},\tilde{\mathbf{z}}^{(2)}]$ (instantiation of Eq.6). Additionally, the graph construction logic is consistent with that of the fine-grained branch.
@@ F.2 Scale analysis
Examples amount 1 2 3 4
@@
Training data ratios ETTh1 ETTm1
@@ F.3 Fully tuning analysis
Method ETTh1 ETTm1
@@ F.4 Comparison between VCA and other LLM-based methods
Methods FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
Methods FSCA VCA S2IP-LLM Time-LLM GPT4TS
Average
SMAPE 11.828 11.889 12.021 12.494 12.690
@@
Methods FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
Methods FSCA VCA S2IP-LLM Time-LLM GPT4TS
@@
Methods LLM-based
@@ F.5 Experimental efficiency analysis
Training Params  Training Params Percentages Training Time for 1 iteration(s) Inference Time for 1 iteration(s)
@@ F.6 Comparison of different prompt types
Variant ETTh1 ETTm1 ETTm1 ETTm2
@@ F.7 Comparison of different network types implementing Context-Alignment
Variant ETTh1 ETTm1 ETTm1 ETTm2
@@ F.8 Comparison with TS Foundation Models
Methods LLM-based Models TS Foundation Models
@@
Methods LLM-based Models TS Foundation Models